Connection from IP address 3.215.180.226
Time: 2021 Mar 09 4:29:42
Found 2591 logbook entries, displaying most recent 100

Logbook Contents (Most Recent Entries)

03 Mar 2021 Wednesday 15:36 [Tom Massey] Alpha source hasd 750 na DC. Swinger Be target 12 channels 2.4 ns width with 12 nA on target!

03 Mar 2021 Wednesday 15:08 [Tom Massey] Alpha Pulsing and bunching. F/4 had 12 na on target. The gamma-width was 12 channels ~2.4 ns.

02 Mar 2021 Tuesday 16:55 [Greg] Sputter source ion gauge was not working. Replaced gauge, found bad gauge. Replaced gauge again, still not working. Controller has failed. Swapped with Lesker controller that was on 45R, leaving 45R with no ion gauge at the moment. Not sure how to adjust the set points on this controller, or what controller to use on 45R. Cooling water pump is also leaking at the shaft seal. Pump is currently shut off, which allows cooling to function with no load, and reduces the volume of the leak. Can turn the pump back on to run, but needs to be swapped for the replacement, and a new seal ordered so that we can have a backup.

26 Feb 2021 Friday 17:13 [Greg] Doug was able to re-manufacture the utility loop water pump and make it run. Installed it, along with 4 new ball valves, and got the utility cooling loop back up and running. Pressures look good, and I bled some air out of the system. Heat exchanger seems to have brought the temperature down to normal. Replacement pump ordered and arrived today to receiving, should be at the lab on Monday. Will work to get it ready to plug in quickly, should this pump fail or develop a leak.

23 Feb 2021 Tuesday 09:28 [Greg] Found water pump making noise this morning, working to investigate. Facilities electricians and controls guys are here to work on the HVAC systems as well.

17 Feb 2021 Wednesday 09:58 [Justin Warren] Came in this morning (2-17-21 @ 9:30) to check the status of my experiment and both alarm panels were in alarm. On the shutdown module, the Cup In light was on, no lights were on main alarm panel. I reset both systems as per the manual and no further alarms occured.

02 Feb 2021 Tuesday 11:27 [Greg] Backfilled Alphatross with Ar, and found that the orifice was not completely plugged. More investigation found some rubidium compound was built up on the canal holder (part 20) and some obstruction of the canal for the RF source (part 24). The source exit flange (part 21) also appeared to have some buildup of rubidium compounds. I was able to clear out the canal, and poke some of the compounds out of the exit flange, but saw some minor sparking, and decided to stop. Source is reassembled now, and pumping down to high vacuum. Can try to start plasma and produce beam, but the source almost certainly need to be torn down and cleaned.

01 Feb 2021 Monday 14:42 [Greg, Don] Still having problems with alphasource. Probe current limited out at 7.5mA several times, forcing a lower probe voltage. Currently shutting down source (2021-02-01 14:30). Once cooled, we will back-fill with Ar and check for blockage of the orifice between the plasma bottle and the charge exchange chamber. Symptoms seem similar to the last blockage, though not identical.

29 Jan 2021 Friday 14:28 [Greg] Monthly safety inspection found that the emergency backup light in the tunnel had failed. I got the spare from Don and will replace it. I'll put together a procurement request for 1 spare to put on the shelf.

27 Jan 2021 Wednesday 08:56 [Greg] Started Alpha source. Left Probe at 6kV, current was around 3mA. Did not start oven yet, but plasma is running well.

26 Jan 2021 Tuesday 17:02 [carter,Greg] Documenting alphatross heater voltages and currents for proper operation. Heater is 120V 200W. At 200W, 120V would be 1.666A. Resistance = 72ohm. Test values using type K thermocouple to confirm temperature probe... 3.1 mV = 95C, 4.0mV = 120C, 5.2 mV = 146C, 7.4 mV = 175C. 75 volts on the supply voltmeter gives 1.04 Amps on the supply ammeter.

25 Jan 2021 Monday 12:47 [Greg, Don] Power outage recovery this morning. Diffusion pump on 15L beamline was not working. Troubleshooting found failed temp switch; replaced with spare from pump in the LTR. Backing pump for 15L beamline also was not operating. Need to replace circuit breaker. Cryo-pump compressor on 45R beamline was low on UHP He. Warmed, refilled, roughed, and restarted compressor. All vacuum systems now operating properly.

23 Jan 2021 Saturday 16:55 [Ingram] LE and HE extensions pumping down.

23 Jan 2021 Saturday 12:33 [Ingram] Sputter source pumping on its diffusion pump. Alphatross now on turbo, I figured out how to start it. Beamline cryopumps cooling down. 15L beamline diffusion pump cold, target chamber diffusion pump working. Ion source cooling on.

23 Jan 2021 Saturday 10:58 [carter] Checked and restored power to electrical and computer systems after power outage. Restarted accelerator logging and alphatross logging.

23 Jan 2021 Saturday 08:18 [Ingram] Power back about 1:30 AM. Came in about 7:30 AM. Reset all alarms. Started all secondary water pumps. Primary water and HVAC all look good. Turned off beam line cryopumps to regenerate. Will come back later to restart cryopumps. Sputter source backing pump and freon compressor on. Alphatross vacuum power on but I don't know how to turn the turbo on.

21 Jan 2021 Thursday 11:23 [carter, Greg] After alphatross repair last Friday we achieved 550 nA DC alphas from the source yesterday 20 Jan 2021. We noticed reduced RF amplitude on the RF meter so made some adjustments in the alligator clip positions on the tank coil and it increased the RF level back close to the previous value which was very close to the red area on the meter. We moved each alligator clip approximately 0.25" to the right facing the source.

21 Jan 2021 Thursday 11:15 [carter] Added alphatross heater voltage plot to the web based monitoring page.

15 Jan 2021 Friday 16:46 [Greg, Don] TL;DR: Alphasource troubles Thursday and Friday. Some problems found and corrected, will try to run the source on Tuesday. Troubleshooting the Alphasource this morning. After overnight vacuum pumping, vacuum levels in the 10^-8 Torr range were achieved. Don and I isolated the probe voltage terminal, and applied voltage to ground. We read high resistance at the 250V and 500V settings, but on the 1000V setting a plasma was induced in the bottle, glowing purple. Based on the wikipedia entry for gas-discharge lamps, it could be argon, nitrogen, or oxygen. This would seem to match with the symptoms that we experienced yesterday working with Tom and Nisha, in that we created a current limiting discharge somewhere around 1000V. This would also seem to indicate that there is something in the plasma bottle other than vacuum or helium, as vacuum shouldn't arc and the helium supply is off, and the plasma is the wrong color. We also noticed that adjusting the mass flow on the helium did not affect the vacuum reading at the ion gauge. Ran roughing vacuum on plasma bottle, and back-filled source with argon. Removed plasma bottle and found no flow of argon. Used 1.58mm rod that Don obtained to clear out the 2mm diameter opening between the plasma bottle and the cesium vapor chamber, at which point argon was able to flow out. Reinstalled the plasma bottle, disconnected argon, rough pumped and allowed the turbo to pump on the source. Pressure currently 2 x 10^-7 Torr and falling. Megger applied to probe voltage contacts, and previously observed discharge is now gone, indicating improved condition. Plan to allow the vacuum to pump until Monday, when we will try to restart the source again.

08 Jan 2021 Friday 14:54 [Greg] Ran windows update on VESDA system, and paused updates until February. Also managed to fix the audio driver, by rolling back to an older version.

06 Jan 2021 Wednesday 08:50 [Greg] Walkthrough of the lab this morning. Backing pump on the swinger seems louder than I remember. This is the pump that we replaced the bearings and coupler on before break. Did not seem excessively warm, nor did I detect excessive vibration. Will monitor.

23 Dec 2020 Wednesday 10:25 [Ingram] Got 1.2 uA in LE cup of 60 keV 4He- and 1.5 uA in HE cup through the machine at 1MV. Analyszer magnet control not working.

23 Dec 2020 Wednesday 07:48 [Ingram] Alphatross Testing Plasma started at about 5 kV on the probe today, yesterday it started when the rf came on. Oven temp is 165 C and I already have 0.2 uA and increasing. Gas still at .35 sccm. Dropping the flow to 0.2 sccm increased the beam to 0.4 uA, oven now 170 C.

22 Dec 2020 Tuesday 17:24 [Ingram] After source shutdown I went out to check on the 4He MFC. It had readings that make sense now. psia 27 degC 22.6 set pt. 0.000 main display 0.000 SCCM He 0.00 ccm 0.000 sccm I think the 4He MFC had to go through some sort of reset for it to work normally. Go figure. David

22 Dec 2020 Tuesday 16:39 [Ingram] Alphatross testing Gas left running all night, no change in pressure still 3E-5 Torr. Shutdown. Went out to check the MFCs. Only strange thing are some negative numbers for flow on the 4He controller. Noted the 3He MFC had 4 psia. Tried to run 3He MFC. at .35 sccm the pressure went to 2E-5 Torr, looks like the same problem. Tried to run down 3He controller and turn off its flow, but, although it was off, it showed a non zero flow of about 0.15 sccm. Used Run All Down and source off and eventually it went to zero. Tried the 4He MFC again and it was working properly, 0.35 sccm 1.6E-6 Torr. Shutdown and reverted to proper control of the source, CB1 plugged back in etc. Started source and it worked as described in the manual. Ran for several hours at 0.2 uA, after 6 hours refocused and re-tuned magnet and got the beam to 0.65 uA. What worked? Cycling the MFCs and the source controls appeared to fix the problem but why? All this time the turbo has been running at 820 Hz. I kept checking just to make sure that the turbo was not the problem with the pressure.

21 Dec 2020 Monday 10:22 [Ingram] Testing Alphatross. The ion gauge pressure is still reading 3E-5 when the gas starts to flow. The pressure is rising to that level within ten seconds of the source being turned on. The pressure is real and can be seen at the inflection magnet when the gate valve to the source is open - same reading 3E-5. Background pressure at the inflection magnet is 3E-6. Running Alphatross from wall power. The MFC is reading correctly and is set for He. The regulator pressure close to the MFC is reading 10 psi, it was reading 16 psi earlier, probably because the source was not taking any gas and the bottle on the wall is set at about 20 psi. Will leave the gas flowing in this mode for 24 hours, all other supplies are off, CB1 is unplugged. Other - the Alphatross ion gauge on the real time data in the control room is difficult to read owing to some high values last week. Other - facilities just came through to turn the steam/hot water back on in the mechanical room.

18 Dec 2020 Friday 16:26 [Ingram] Shutdown source. The pressure dropped imediately to 2E-7 Torr. It had been falling very slowly but was still 3E-5 Torr before shutdown.

18 Dec 2020 Friday 13:18 [Ingram] Starting Alphatross. IG pressure went to 3.3E-5 Torr when plasma started. Plasma started when RF came on. Will run only RF, magnet and probe until pressure drops to low sixes.

14 Dec 2020 Monday 16:01 [Greg, Devon] TL;DR: Still no vacuum on swinger, turbo might be broken. Swinger work today: replaced ion gauge with the last spare from the shelf. Reinstalled the backing pump. Turbo pump controller is not working, lights just blink, nothing on the LCD, no lights on the turbo pump. Nothing illuminating in the manual, will investigate further as time permits.

11 Dec 2020 Friday 15:20 [carter] Logbook test.

11 Dec 2020 Friday 12:21 [carter] Again, as on 14 Sept 2020, the capacitor in LE tube vacuum gauge was shorted and was replaced. 0.5 uF 4000 VDC. Last spare on the shelf. Used the one salvaged from spare gauge power supply unit that was removed from old duoplasmatron source. Probably only available now from ebay. Checked the supply voltage and it is within ratings, 3600 VDC.

11 Dec 2020 Friday 12:10 [Greg, Devon] Replaced belts on large roughing pump in the vault. Last belt replacement was 2016. Belts were slipping from oil and ragged. Oil level in the sight glass was low, need to think about an oil change soon.

11 Dec 2020 Friday 09:44 [Greg, Devon] Replaced ion gauge on 65R. Roughed out chamber and beam line.

10 Dec 2020 Thursday 12:10 [Devon Jacobs] Came in this morning to find a phase failure alarm. Found many of the usual suspects, ion gauges, gate valves, in an off state. Of particular note was the alphatross vacuum pumping system. We didn't discover the source vacuum was out until much later. The only indication of such a failure is visual inspection of the vacuum control module in the source area, or if we had thought to look at the software which would have been displaying whatever it would with the ion gauge off. Going forward, should consider tapping into vacuum control circuitry to establish an alarm or other kind of indication for a failure. In other news, facility shutdown is progressing. We've also found the process cooling water emergency dump valve to be inop. Solution in work...

04 Dec 2020 Friday 11:34 [carter, leblanc] Compressor blew fuses again. A new set of contacts has been installed. Also a missing bolt was replaced on the solenoid actuator. Hopefully one or the other was the cause of blowing fuses.

02 Dec 2020 Wednesday 16:08 [carter] Test of addition of Greg LeBlanc to logbook "mail to all operators" function.

02 Dec 2020 Wednesday 16:07 [carter] Test of logbook "mail to Greg LeBlanc" addition.

18 Nov 2020 Wednesday 14:37 [carter] The main air compressor went out again today during the run shutting down the machine. Found one blown fuse and one thermal overload tripped again, same as before. This time I have burnished the contacts as this is a possible cause. I will also check on the availability of a new set of contacts for this contactor. Part number is Square D TA81. This set was replaced for the same symptom April 2006.

18 Nov 2020 Wednesday 11:41 [carter] Currently there are software imposed limits for the alphatross controls of the following, Mass Flow for HE3 and HE4 of 0.18 to 0.50 SCCM and also a software limit on the oven voltage of 90 volts max.

18 Nov 2020 Wednesday 11:25 [Tom Massey] Alphatross start up normal. Starting current 1.1-1.2 microamper. This was slighly unstable. Dr. Ingram saw this behavior also but this eventually stablized at the higher value. His test showed that the source mass flow can be reduced to ~0.18 CCM with improved current output. We currently have 200 nA on target with no current on the collimator at the 65R HeBGB target position.

13 Nov 2020 Friday 11:23 [Devon] The beam right object slit is repaired. Took the micrometer from the Y-up slit. Did not have to break vacuum. The failed beam right slit was not at or beyond the zero point, it seemed to be sitting at around the 0.03" open point. But that is only an estimate as to how far open it was. New Beam right zero set when slits were in contact with beam left at zero. The Y-up slit is trapped in place at > 0.4" setting with a hose clamp to prevent the spring from pushing the slit out further.

11 Nov 2020 Wednesday 15:27 [Tom Massey] The neutron sources have been returned to the source room.

08 Nov 2020 Sunday 10:38 [Tom Massey] The beam right object slit no longer moves. Trying to turn by hand is not possible. The shaft does not move at all. We have been switching the object and image slits between 0.030 and 0.200. This is done to determine the energy by the magnet and calibrate the nmr at the smaller setting of 0.030. The slits are then opened to get enough beam to obtain enough counts above background.

07 Nov 2020 Saturday 07:54 [Tom Massey] The alphatross has 1.1 microamps this morning.

06 Nov 2020 Friday 09:14 [Tom Massey] The swinger turbo pump's roughing pump was making loud noises this morning. The turbo pump has been turned off.

04 Nov 2020 Wednesday 16:16 [carter. shafer] At 11:30 PM last evening the main air compressor stopped. Upon inspection found and repaired the following, blown fuse, thermal protector tripped in motor starter relay box, conduit feed termination had worked loose and needed re-connected at motor and a long bolt (1 of 4) that holds the motor bearing housings together had worked loose and needed replacing. Doug had a new piece (1/4 in. all-thread) in the shop to replace the bolt. After the repairs were made the motor started and the compressor is working again.

03 Nov 2020 Tuesday 23:30 [Doug Soltesz] Tonight at 11:20 pm, the "water hot or air low" alarm sounded and the grid voltage quickly increased to around -15 V. The accelerator was immediately shut off and the magnets turned down. We contacted Devon, who said it was probably the air compressor, and he will be in to fix it on Thursday.

03 Nov 2020 Tuesday 23:30 [Doug Soltesz] Tonight at 11:20 pm, the "water hot or air low" alarm sounded and the grid voltage quickly increased to around -15 V. The accelerator was immediately shut off and the magnets turned down. We contacted Devon, who said it was probably the air compressor, and he will be in to fix it on Thursday.

02 Nov 2020 Monday 08:09 [phase failure detected] The phase failure detected alarm on the E2 panel was on this morning. I checked all vac systems and everything seems to be operating normally. Best guess is there was a very short power outage that set off the alarm but did not hurt anything. I will proceed to turning on the source unless I hear otherwise. -KB

30 Oct 2020 Friday 13:31 [carter] Emergency generator oil changed by facilities.

28 Oct 2020 Wednesday 17:40 [carter] The fictitious chain stretch alarm was caused by sparking at the alphatross source. The cause was a diagnostics induced disconnected plasma box fan cable that was not plugged in and hanging from the source platforming sparking to ground. Initial restart of the source produced 450 nA of beam within 15 minutes.

28 Oct 2020 Wednesday 12:48 [Tom Massey] The chain stretch alarm is going off. This happened when the source was being brought up. This came on when the bias voltage was being raised. The alarm was gone when the bias voltage on the alphatross was reduced. When the extractor voltage reached above ~5 kV the alarm would be set off. Sparking???

23 Oct 2020 Friday 15:44 [Tom Massey] The accelerator was conditioned up to 4.5 MV. We are currently running at 3.371 MV.

15 Oct 2020 Thursday 17:15 [Devon] Alphatross update: All 10 g of Rb loaded without issue. Flowed He thru the source over lunch. Then went for operation. Plasma struck right away but vacuum reading was higher than expected as Probe voltage came up. Let it run a couple hours as suggested in the manual, vacuum came down to expected readings. Our instructions have been changed to say the Probe runs at 6 KV instead of 6.5 KV. Not sure how we got to 6.5, but the manual and other labs use 6. Moved on to oven heating and applied voltages. The focus is a key player in tuning now, which we didnít see when the focus element was shorted by Rb. Initially the beam output was less than 100 nA and the gas flow was higher than expected. As time went on a lower gas flow started giving higher beam output, now closer to the expected .2 sccm. I recall seeing this the last time the bottle was changed. The Rb oven behavior is different now. I had used a bit of thermal contact grease on the TC last time. This time I packed the connection heavier with grease. The oven was set to 87.5 V, what had become our standard and gave us a temperature reading in the upper 180s C. We had seen instability appear if the temp reached 190 C around 90 V. After some time of running today it was noticed that the oven temp had already reached 191 C at 87.5 V. The beam had also come up to 250 nA. Lowered the oven volts to get temp back down and then started nursing it back up. We are currently at 185 C at 77.5 V and getting 200 nA. This makes me think that we were not getting good TC readings on the last build and actually over driving the oven. Something had to make the Rb mess. Iím going to leave the source running overnight as it is right now. Perhaps the beam will improve by morning. If or if not, there should be some exploration of new operating parameters on the oven. No idea what the upper limit is at this point. If somebody wishes to explore this tomorrow then great. If not, please shut down the source. Also, the big einzel lens pretty much does nothing. If you raise it high enough it defocuses the beam a little. Another item for further contemplationÖ

08 Oct 2020 Thursday 11:39 [Devon] Alphatross update: It took many iterations of wiping with Kimwipes and water, acetone, alcohol, and heat gun to get the voltage gap insulators to measure >4000 on the MegOhm meter. The extractor gaps were worse than the acceleration tube. The extractor/lens core components all had to be bead blasted, the Al items were fairly pitted by Rb. The source is back together and under vacuum. New plasma bottle and canal installed. Nothing remarkable about right/left alignment on the probe electrode, but it was observed that pushing on the gas line can move the bottle right/left some. The Ta canal tube had eroded on the exit end, almost thru to the Al base plate. Seems that this is the heaviest wear item along with the Al canal holder button. The new RGA is installed and running. Initial analysis of just vacuum showed no indication of leaks, typical High Vac spectrum according to DCI, although there is a hydrocarbon peak, which is odd since this is a dry pumped system. Now flowing He for clean up and watching the RGA. Masses 28 and 32 come on pretty strong when the gas flowsÖwill continue to monitor this.

29 Sep 2020 Tuesday 11:59 [Devon] Alphatross update: The CEC is clean and re-assembled. Removed the extractor/lens assembly. Puddle of Rb behind (downstream) the extractor electrode opening. Good look down the acceleration tube, there is a big collimator plate in front of the steerers, very large blue stain on the Ta plate which is centered. Very large gold stain on the Ta plate which is off to the right and low. Not sure what to make of the size, color, or location of these stains. Beginning the tear down and cleaning of the extractor/lens...

28 Sep 2020 Monday 15:04 [Tom Massey] All of the neutron souces have been returned to the source room.

25 Sep 2020 Friday 15:38 [Devon] Alphatross notes on disassembly: Iíve posted a drawing of the source on the handbook to help with any references below (why does the Comdoc always make a crooked scan?). Rb neutralized without much fanfare. The greatest reactions were in the vault during disassembly, but not enough to register on the VESDA fire alarm. Care and patience must be taken when removing the oven, the wire wick has lots of Rb that gets re-exposed as the wick is moved. The water tub outside was unremarkable. Not much Rb left in the oven, guessing maybe 20% or less of the initial charge. The plasma bottle has milky white coating between the RF bands which Iím thinking is an expected result. The insulator and exit canal seem to be in quite good shape. Still plan to use a fresh bottle and clean/align the probe assembly (gas end, items 28-36). The Al canal holder, item 20 (exit aperture) was again eroded off axis to the right. This is why I want to take a good look at the probe end of the bottle. The holder was not very caked with Rb. Next we get to the charge exchange chamber (CEC) entry cooling baffle (item 48), its aperture was well occluded, less than a quarter of the original diameter (6.4mm) by eye, and the remaining opening was off axis right. The CEC itself (all of item 17) didnít seem too bad, nothing like the ANSTO pictures on SNEAP. The entry aperture was about half occluded (nominally 4.8mm), and there was some Rb puddling at the base of the flange where it meets the entry cooling baffle. The CEC exit was interesting. The Source Exit Spacer (Stainless steel, item 3) has a much larger hole (12.75mm) and conically expands after the hole until it meets the titanium exit aperture (item 4) which has the same size hole. The idea is that Rb should condense on the exit aperture and flow back to the oven thru the drain hole in the bottom of the conical spacer as part of the Rb cycle. Some Rb had wicked into the contact gap between aperture plate and spacer, but the drain was open, so one would assume the cycle is working. Any Rb that escapes the exit aperture is confronted by the exit cooling baffle (item 52, 19mm diameter) and the large space between the source base plate (item 1) and the baffle plate. This space had a fair amount of Rb pooled in the bottom, the greatest concentration other than the oven. Rb that collects in this space is gone from the cycle. The next components of the source are the extractor/lens (focus) assembly followed by the acceleration tube (bias). Gap resistance was checked on these devices while the source was still under vacuum. The 300 MOhm resistor gaps on the acceleration tube measured 240 MOhm consistently. The 150 MOhm resistor gaps on the extractor measured 100 MOhm consistently. Didnít think I needed to check continuity btween the lens electrode and the adjacent extractor gaps. We re-checked resistivity after the source was vented to air. The acceleration gaps had now dropped to 80 MOhms and decreased continuously towards the CEC. The first extractor gap had dropped to 1 MOhm and the others got less moving towards the CEC. Checked the lens to the adjacent extractor plane closest to the CEC, 300 KOhms; adjacent plane further from the CEC, 600KOhms. Both of those measurements should have been infinite since the lens element is isolated. Just for fun, removed one 150 MOhm resistor, measured 149 MOhm. So, lots more cleaning and disassembly ahead of us.

25 Sep 2020 Friday 15:12 [Tom Massey] The Pu(Be) and 252Cf sources have been placed in the back of the tunnel. Kristyn is doing calibrations and needs a low neutron background.

24 Sep 2020 Thursday 13:59 [Devon] Alphatross notes prior to disassembly: 881.6 hours. Generally suffering from low beam output and instability. After the last power outage and Ar backfill the source sat under vacuum for several weeks of sputter source runs. Upon start up there was immediate plasma ignition with a pinkish hue. Source was left running in hopes of clean up, the pinkish hue faded to aqua quickly. He- was initially on the order of 100-300 nA in the LE cup with instabilities on the order of 5-10 Hz. After several days the output had increased to around 500 nA but the source had to be shutdown for sputter pellet change. On restart the plasma would not ignite, had to raise the gas to .5 sccm and raise the probe to 5 KV to get ignition. After another 4 days the current had increased to 800 nA with instability on the order of 1-2 Hz. During RBS work on day 5 the Probe was increased from 6 to 6.5 KV which gave an increase in beam reaching the RBS chamber. But from that point forward the beam suffered a steady downward slope for a few days until the output had reached 100 nA again, and the magnitude of the instabilities had increased, but not the frequency. Attempts at tuning couldnít produce any more than 150 nA at the peak of the instability cycle. The BPM showed a reasonable looking profile at times, but it was also swamped with baseline noise, the effect seen when ions strike the housing. In some cases the baseline noise would persist for many seconds after a voltage change in the system. Nothing left to do but tear down the sourceÖ

24 Sep 2020 Thursday 12:05 [carter] OIT has returned the portable phone for the vault claiming is is OK. I have returned it to its normal place over the swinger scope inside the vault door. We will see.

21 Sep 2020 Monday 14:53 [Gula Hamad] Note on BCI Calibration: BCI test was done using Both ranges 1E-06 and 1E-07 and we got the same result.

21 Sep 2020 Monday 14:33 [Gula Hamad] BCI Calibration: We did test the BCI by sending a known current and verifying the BCI reading. We calculated the measured charge from BCi = 5.97E-04 [C] and the Applied current = 6.00E-04 [C]. The ratio between charge from BCi to the Applied current = 0.9955

19 Sep 2020 Saturday 13:46 [Ingram] Alphatross note: I saw that the probe was at 6.0 kV. I decided to increase it to 6.5 kV. The beam on target ( I am taking data) increased, it doubled from 10 nA to 20 nA.

18 Sep 2020 Friday 12:59 [carter] With reference to the buncher no output issue, I decided to double check that the tank switch was indeed in position 2, rotated the switch completely around and back to position 2, the position we wanted for alphas. Re checked the operation. Saw a glimmer of 5 MHZ on the output and tuned the output tune control many turns and found the resonance, so it appears that either the tube selector switch was not completely in position or possibly that we don't frequently run in position 2 and it needs to be re-tuned significantly to resonate. Anyway, it appears that the buncher is again OK and tuned for the 16 cm bunch tube. Pulsed beam should now again be possible.

18 Sep 2020 Friday 11:15 [carter] The portable phone that was located in the vault has been picked up by OIT for diagnostics and repair. There is no phone at that location now so use the portable from the control room if phone service in that area is needed until further notice.

18 Sep 2020 Friday 11:10 [carter] The buncher has developed a failure overnight. Restricted access diagnosis is a defective output tube screen power supply. DC operations can continue with the machine as scheduled. Maintenance on the buncher is scheduled for Sept 30 unless an opportunity presents itself and the sources are down before that date.

17 Sep 2020 Thursday 09:12 [Devon] Magnet Primary pump running. Resume schedule.

16 Sep 2020 Wednesday 15:49 [Devon] A viable bearing assembly for the magnet pump was found in the shop. Still has bearings, yay. Hope to have the existing pump up and running tomorrow, for as long as the motor lasts.

16 Sep 2020 Wednesday 15:45 [Devon] Additional videos have been added to the bottom of the handbook page. There is a video of navigating the control room camera station and a startup and shutdown video for each source. Enjoy.

15 Sep 2020 Tuesday 15:55 [Devon] The Magnet Primary cooling pump failed and is unrepairable. Attempted to install a new shaft seal and coupling but wear and pitting were too bad to get a seal. This was only going to be temporary anyway due to severe slop in motor bearings. Checking on price and lead time for a replacement...

15 Sep 2020 Tuesday 11:29 [Devon] Magnet Primary water pump failed. Attempting rebuild...

14 Sep 2020 Monday 16:22 [carter] Capacitor in LE tube vacuum gauge was shorting and was replaced. 0.5 uF 4000 VDC. Last spare on the shelf but have a complete spare gauge power supply unit that was removed from old duoplasmatron source as a supply of spares.

11 Sep 2020 Friday 11:51 [carter] Analyzing magnet power supply would not run up. Problem traced to the collector mount screw on Q-7 that had loosened, probably because of thermal cycling. Will add star locking washers to these screws to help prevent this.

05 Sep 2020 Saturday 14:39 [Ingram] pellet changed to 12C. Left source gate valve closed.

04 Sep 2020 Friday 16:43 [Devon] The zero gauge reading was on the suction side of the comfort chilled water pump. The make up regulator was stuck and there was a leak in the Trane chiller. The regulator set point was exorcised and is working again. The leak was a small valve with a cracked body feeding an air bleeder inside the chiller. Closing the valve stopped the leak for now. There is a work order to replace the cracked valve.

04 Sep 2020 Friday 11:23 [Ingram] Beam instability back so that I have switched to GVM control. This coincided with instabilty in the extractor current in the alphtross.

04 Sep 2020 Friday 10:40 [Ingram] Water in the mechcanical room above the control room is coming from three sources. Water from the HVAC air handler, as usual. Water from a roof leak probably from the rain yeasterday giving a pool of water around the water heater. This is dripping very slowly through to the control room and onto the table close to my laptop, that is why I noticed. The third source is in the inner room around the water pump. The blue cylinder connected to the water pump is sweating alot. Also one of the water gauges connected to the HVAC pump is reading zero.

04 Sep 2020 Friday 10:36 [Ingram] Alphatross producing about 0.7 uA LE cup. Stable. Running 10 nA on target and taking RBS and PIXE data. Things of note about the alphatross, to me, Einzel lens, Focus Voltage and both steerers are at zero volts. They either have no effect on the beam or in the case of the steerers, reduce the current. No obvious reason why the beam current has grown over time or why the beam is now more stable.

03 Sep 2020 Thursday 12:04 [carter] Replaced V6 6AH6 120 Hz drive tube in the NMR to restore the full horizontal sweep in the 120 Hz position. Adjusted the sweep gain to match the 60 Hz.

02 Sep 2020 Wednesday 15:47 [carter] Created a wtchdog timer for the accelerator logging system. If the logging fails to update the tandem.time file over a 5 minute period, the system will automatically restart the acelerator logging program.

02 Sep 2020 Wednesday 08:20 [carter] Looks like accelerator logging system stopped at 2:23:54 AM this morning, possibly power glitch that shut off power to sputter source HV platform? Anyway, restarted the logging system at 8:12 this morning. Looks like it is working OK now.

31 Aug 2020 Monday 08:39 [Devon] The front door is being replaced this morning.

28 Aug 2020 Friday 12:18 [carter] The alphatross computer has completed the new disc image. The new image was installed as the current system disk, the old one retired as a backup disc. The new system has been updated to the current centos7 version 7.8.2003. Tandemwatch and inflection controls restored to normal.

27 Aug 2020 Thursday 12:41 [carter] I have begun an image backup of the alphatross computer prior to applying system updates for Centos7. The inflection magnet control is available on the tandem computer and the tandemwatch (downcharge monitor program) is displaying on the congrol room ddaq computer. I will restore the alphatross computer to normal operation Monday AM.

24 Aug 2020 Monday 15:20 [Zach Meisel] Operators needed (week of September 7): Greetings operators, Alexander will be running 12C on Mg targets in the charged-particle time-of-flight chamber on R65 to study the statistical properties of nuclides in this region. He would like operator help to do this. If possible, we will try for overnight runs. These runs will mostly involve sitting at a single energy and letting the data come in. Please fill-out this Doodle poll to indicate your availability: https://doodle.com/poll/uwrhze3nnkada3sb. Thanks for your help! -Zach

24 Aug 2020 Monday 12:52 [Cody Parker] Came in to stop the overnight TexAT calibration and the SF6 alarm was going off. I lowered the set point slightly and the alarm stopped. I checked the tank pressure and temperature, and both have dropped some. I talked to Don on his way to the LTR and he will monitor the plots.

23 Aug 2020 Sunday 20:25 [Cody Parker/Jack Bishop] We have completed the 12C(n,n2) and 16O(n,a) runs using TexAT! Thank you to everyone for all of your help! We are working on a calibration source run with TexAT overnight, and plan to start the ~week-long process of disassembly and packing tomorrow. The D2 gas cell is still pressurized and the water cooling is on. All other accelerator-related equipment and the sputter source have been shut down.

23 Aug 2020 Sunday 00:20 [Devon] VESDA minor smoke event in the small target room. Definite smell of something burned, but no obvious source, detection returned to baseline. Will have to be on the lookout for something not working in there. Came in to reset the fire alarm, experiment resuming.

12 Aug 2020 Wednesday 14:07 [carter] Trane replaced other temperature sensor today and both circuits are said to be OK and running now.

11 Aug 2020 Tuesday 16:14 [Devon] The gas is in and drying. The accelerator is ready to run tomorrow as long as the AC stays on.

11 Aug 2020 Tuesday 11:49 [Devon] Chiller update: A factory tech was here this morning. He happened to have a water temperature sensor. It has been installed and we should be back to half a chiller. They identified another bad sensor that they did not have on hand. Will be back tomorrow late morning or early afternoon to install the other sensor. People will have to be on the look out for the factory tech, and pause the run or survey the roof, and let him in the roof gate to make the repairs. I have started to evacuate the accelerator tank. Gas transfer will be done by 5 pm if the door seal doesn't leak.

11 Aug 2020 Tuesday 07:17 [Devon] The Facilities work center page is still not functioning on VPN. The issue was reported to OIT yesterday morning. There was an urgent work order for the chiller in early July that was never addressed. I can't obtain that work order number until I get to campus.

11 Aug 2020 Tuesday 07:13 [Devon] So the chiller was completely dead by around 7 pm last night. First person in the lab today needs to call Facilities and report a chiller emergency, and stand by to let in the technician. I won't be there until 9:45ish. From what I discussed with the technician yesterday, a reliable cure for this chiller is not certain and certainly not timely. The thing is a lemon. Management from both organizations need to discuss a timeline, consider installation of a portable, temporary chiller, and discuss the disposition of our out-of-state visiting researchers.

10 Aug 2020 Monday 14:48 [Devon] Facilities Management was here most of the day. Both halves of the chiller are still running at this moment. There is still a problem such that the measured water temperature in the unit and the actual water temperature are different. They will order new sensors. But its running. Plan is to transfer gas tomorrow after operator class.

10 Aug 2020 Monday 11:53 [Devon] Tank Opening: 7173 hours. We observed a sudden onset of noise in the LE Tube downcharge while running at 2.7 MV. We entered the tank expecting to find an open resistor or broken connecting spring. We found neither. We started the process of looking back at logged data. The onset of the noise was related to the tube entrance lens being driven to a voltage above the terminal induced voltage of the downcharge on plane 4 where the lens is. We investigated the performance of just the lens with vacuum in the tube but 1 atm air in the tank. The lens voltage and current were stable and giving the expected values. There is a spark gap on the lens feed path. The points of the gap were dirty and sparked at around 8 KV. We insulated the gap with Teflon sheet and ran at around 15 KV without instability, confirming that the supply and wiring seemed to be ok. A review of the prior month data showed that the noise had been intermittently present when the lens was driven above the induced voltage, or when the lens had been left on while the accelerator is off (donít leave electrostatic devices on!). Given the compulsion to do SOMETHING while the tank was open, decided to clean the spark gap with sandpaper and kimwipes. Nothing else to do with respect to this problem. We will have to monitor the noise behavior going forward. It had also been a year since the bearing disaster. A year is the target timeframe for greasing bearings, so that was done. No other maintenance due to schedule pressure.

10 Aug 2020 Monday 06:34 [carter] Same thing for me, says not on VPN but I am also.

10 Aug 2020 Monday 06:17 [Devon] VPN test. Facilities web page says I'm not on VPN, page won't work. VPN says connected, this page works...

07 Aug 2020 Friday 09:48 [Devon] During the monthly safety inspection, found the vault phone to be severely locked up. Had to remove the battery. Seems to be OK now. Radiation?

05 Aug 2020 Wednesday 16:26 [Tom Massey] We were running at 2.5 MV on the terminal today. Devon noticed that the LE current has been jumpy. This did not change when the LE cup was put in. These excursions seemed to reduce in size with time. The changes were most visible on the Tandem watch screen.

Connection from IP address 3.215.180.226
Time: 2021 Mar 09 4:29:42
Found 2591 logbook entries, displaying most recent 100

Logbook Contents (Most Recent Entries)

03 Mar 2021 Wednesday 15:36 [Tom Massey] Alpha source hasd 750 na DC. Swinger Be target 12 channels 2.4 ns width with 12 nA on target!

03 Mar 2021 Wednesday 15:08 [Tom Massey] Alpha Pulsing and bunching. F/4 had 12 na on target. The gamma-width was 12 channels ~2.4 ns.

02 Mar 2021 Tuesday 16:55 [Greg] Sputter source ion gauge was not working. Replaced gauge, found bad gauge. Replaced gauge again, still not working. Controller has failed. Swapped with Lesker controller that was on 45R, leaving 45R with no ion gauge at the moment. Not sure how to adjust the set points on this controller, or what controller to use on 45R. Cooling water pump is also leaking at the shaft seal. Pump is currently shut off, which allows cooling to function with no load, and reduces the volume of the leak. Can turn the pump back on to run, but needs to be swapped for the replacement, and a new seal ordered so that we can have a backup.

26 Feb 2021 Friday 17:13 [Greg] Doug was able to re-manufacture the utility loop water pump and make it run. Installed it, along with 4 new ball valves, and got the utility cooling loop back up and running. Pressures look good, and I bled some air out of the system. Heat exchanger seems to have brought the temperature down to normal. Replacement pump ordered and arrived today to receiving, should be at the lab on Monday. Will work to get it ready to plug in quickly, should this pump fail or develop a leak.

23 Feb 2021 Tuesday 09:28 [Greg] Found water pump making noise this morning, working to investigate. Facilities electricians and controls guys are here to work on the HVAC systems as well.

17 Feb 2021 Wednesday 09:58 [Justin Warren] Came in this morning (2-17-21 @ 9:30) to check the status of my experiment and both alarm panels were in alarm. On the shutdown module, the Cup In light was on, no lights were on main alarm panel. I reset both systems as per the manual and no further alarms occured.

02 Feb 2021 Tuesday 11:27 [Greg] Backfilled Alphatross with Ar, and found that the orifice was not completely plugged. More investigation found some rubidium compound was built up on the canal holder (part 20) and some obstruction of the canal for the RF source (part 24). The source exit flange (part 21) also appeared to have some buildup of rubidium compounds. I was able to clear out the canal, and poke some of the compounds out of the exit flange, but saw some minor sparking, and decided to stop. Source is reassembled now, and pumping down to high vacuum. Can try to start plasma and produce beam, but the source almost certainly need to be torn down and cleaned.

01 Feb 2021 Monday 14:42 [Greg, Don] Still having problems with alphasource. Probe current limited out at 7.5mA several times, forcing a lower probe voltage. Currently shutting down source (2021-02-01 14:30). Once cooled, we will back-fill with Ar and check for blockage of the orifice between the plasma bottle and the charge exchange chamber. Symptoms seem similar to the last blockage, though not identical.

29 Jan 2021 Friday 14:28 [Greg] Monthly safety inspection found that the emergency backup light in the tunnel had failed. I got the spare from Don and will replace it. I'll put together a procurement request for 1 spare to put on the shelf.

27 Jan 2021 Wednesday 08:56 [Greg] Started Alpha source. Left Probe at 6kV, current was around 3mA. Did not start oven yet, but plasma is running well.

26 Jan 2021 Tuesday 17:02 [carter,Greg] Documenting alphatross heater voltages and currents for proper operation. Heater is 120V 200W. At 200W, 120V would be 1.666A. Resistance = 72ohm. Test values using type K thermocouple to confirm temperature probe... 3.1 mV = 95C, 4.0mV = 120C, 5.2 mV = 146C, 7.4 mV = 175C. 75 volts on the supply voltmeter gives 1.04 Amps on the supply ammeter.

25 Jan 2021 Monday 12:47 [Greg, Don] Power outage recovery this morning. Diffusion pump on 15L beamline was not working. Troubleshooting found failed temp switch; replaced with spare from pump in the LTR. Backing pump for 15L beamline also was not operating. Need to replace circuit breaker. Cryo-pump compressor on 45R beamline was low on UHP He. Warmed, refilled, roughed, and restarted compressor. All vacuum systems now operating properly.

23 Jan 2021 Saturday 16:55 [Ingram] LE and HE extensions pumping down.

23 Jan 2021 Saturday 12:33 [Ingram] Sputter source pumping on its diffusion pump. Alphatross now on turbo, I figured out how to start it. Beamline cryopumps cooling down. 15L beamline diffusion pump cold, target chamber diffusion pump working. Ion source cooling on.

23 Jan 2021 Saturday 10:58 [carter] Checked and restored power to electrical and computer systems after power outage. Restarted accelerator logging and alphatross logging.

23 Jan 2021 Saturday 08:18 [Ingram] Power back about 1:30 AM. Came in about 7:30 AM. Reset all alarms. Started all secondary water pumps. Primary water and HVAC all look good. Turned off beam line cryopumps to regenerate. Will come back later to restart cryopumps. Sputter source backing pump and freon compressor on. Alphatross vacuum power on but I don't know how to turn the turbo on.

21 Jan 2021 Thursday 11:23 [carter, Greg] After alphatross repair last Friday we achieved 550 nA DC alphas from the source yesterday 20 Jan 2021. We noticed reduced RF amplitude on the RF meter so made some adjustments in the alligator clip positions on the tank coil and it increased the RF level back close to the previous value which was very close to the red area on the meter. We moved each alligator clip approximately 0.25" to the right facing the source.

21 Jan 2021 Thursday 11:15 [carter] Added alphatross heater voltage plot to the web based monitoring page.

15 Jan 2021 Friday 16:46 [Greg, Don] TL;DR: Alphasource troubles Thursday and Friday. Some problems found and corrected, will try to run the source on Tuesday. Troubleshooting the Alphasource this morning. After overnight vacuum pumping, vacuum levels in the 10^-8 Torr range were achieved. Don and I isolated the probe voltage terminal, and applied voltage to ground. We read high resistance at the 250V and 500V settings, but on the 1000V setting a plasma was induced in the bottle, glowing purple. Based on the wikipedia entry for gas-discharge lamps, it could be argon, nitrogen, or oxygen. This would seem to match with the symptoms that we experienced yesterday working with Tom and Nisha, in that we created a current limiting discharge somewhere around 1000V. This would also seem to indicate that there is something in the plasma bottle other than vacuum or helium, as vacuum shouldn't arc and the helium supply is off, and the plasma is the wrong color. We also noticed that adjusting the mass flow on the helium did not affect the vacuum reading at the ion gauge. Ran roughing vacuum on plasma bottle, and back-filled source with argon. Removed plasma bottle and found no flow of argon. Used 1.58mm rod that Don obtained to clear out the 2mm diameter opening between the plasma bottle and the cesium vapor chamber, at which point argon was able to flow out. Reinstalled the plasma bottle, disconnected argon, rough pumped and allowed the turbo to pump on the source. Pressure currently 2 x 10^-7 Torr and falling. Megger applied to probe voltage contacts, and previously observed discharge is now gone, indicating improved condition. Plan to allow the vacuum to pump until Monday, when we will try to restart the source again.

08 Jan 2021 Friday 14:54 [Greg] Ran windows update on VESDA system, and paused updates until February. Also managed to fix the audio driver, by rolling back to an older version.

06 Jan 2021 Wednesday 08:50 [Greg] Walkthrough of the lab this morning. Backing pump on the swinger seems louder than I remember. This is the pump that we replaced the bearings and coupler on before break. Did not seem excessively warm, nor did I detect excessive vibration. Will monitor.

23 Dec 2020 Wednesday 10:25 [Ingram] Got 1.2 uA in LE cup of 60 keV 4He- and 1.5 uA in HE cup through the machine at 1MV. Analyszer magnet control not working.

23 Dec 2020 Wednesday 07:48 [Ingram] Alphatross Testing Plasma started at about 5 kV on the probe today, yesterday it started when the rf came on. Oven temp is 165 C and I already have 0.2 uA and increasing. Gas still at .35 sccm. Dropping the flow to 0.2 sccm increased the beam to 0.4 uA, oven now 170 C.

22 Dec 2020 Tuesday 17:24 [Ingram] After source shutdown I went out to check on the 4He MFC. It had readings that make sense now. psia 27 degC 22.6 set pt. 0.000 main display 0.000 SCCM He 0.00 ccm 0.000 sccm I think the 4He MFC had to go through some sort of reset for it to work normally. Go figure. David

22 Dec 2020 Tuesday 16:39 [Ingram] Alphatross testing Gas left running all night, no change in pressure still 3E-5 Torr. Shutdown. Went out to check the MFCs. Only strange thing are some negative numbers for flow on the 4He controller. Noted the 3He MFC had 4 psia. Tried to run 3He MFC. at .35 sccm the pressure went to 2E-5 Torr, looks like the same problem. Tried to run down 3He controller and turn off its flow, but, although it was off, it showed a non zero flow of about 0.15 sccm. Used Run All Down and source off and eventually it went to zero. Tried the 4He MFC again and it was working properly, 0.35 sccm 1.6E-6 Torr. Shutdown and reverted to proper control of the source, CB1 plugged back in etc. Started source and it worked as described in the manual. Ran for several hours at 0.2 uA, after 6 hours refocused and re-tuned magnet and got the beam to 0.65 uA. What worked? Cycling the MFCs and the source controls appeared to fix the problem but why? All this time the turbo has been running at 820 Hz. I kept checking just to make sure that the turbo was not the problem with the pressure.

21 Dec 2020 Monday 10:22 [Ingram] Testing Alphatross. The ion gauge pressure is still reading 3E-5 when the gas starts to flow. The pressure is rising to that level within ten seconds of the source being turned on. The pressure is real and can be seen at the inflection magnet when the gate valve to the source is open - same reading 3E-5. Background pressure at the inflection magnet is 3E-6. Running Alphatross from wall power. The MFC is reading correctly and is set for He. The regulator pressure close to the MFC is reading 10 psi, it was reading 16 psi earlier, probably because the source was not taking any gas and the bottle on the wall is set at about 20 psi. Will leave the gas flowing in this mode for 24 hours, all other supplies are off, CB1 is unplugged. Other - the Alphatross ion gauge on the real time data in the control room is difficult to read owing to some high values last week. Other - facilities just came through to turn the steam/hot water back on in the mechanical room.

18 Dec 2020 Friday 16:26 [Ingram] Shutdown source. The pressure dropped imediately to 2E-7 Torr. It had been falling very slowly but was still 3E-5 Torr before shutdown.

18 Dec 2020 Friday 13:18 [Ingram] Starting Alphatross. IG pressure went to 3.3E-5 Torr when plasma started. Plasma started when RF came on. Will run only RF, magnet and probe until pressure drops to low sixes.

14 Dec 2020 Monday 16:01 [Greg, Devon] TL;DR: Still no vacuum on swinger, turbo might be broken. Swinger work today: replaced ion gauge with the last spare from the shelf. Reinstalled the backing pump. Turbo pump controller is not working, lights just blink, nothing on the LCD, no lights on the turbo pump. Nothing illuminating in the manual, will investigate further as time permits.

11 Dec 2020 Friday 15:20 [carter] Logbook test.

11 Dec 2020 Friday 12:21 [carter] Again, as on 14 Sept 2020, the capacitor in LE tube vacuum gauge was shorted and was replaced. 0.5 uF 4000 VDC. Last spare on the shelf. Used the one salvaged from spare gauge power supply unit that was removed from old duoplasmatron source. Probably only available now from ebay. Checked the supply voltage and it is within ratings, 3600 VDC.

11 Dec 2020 Friday 12:10 [Greg, Devon] Replaced belts on large roughing pump in the vault. Last belt replacement was 2016. Belts were slipping from oil and ragged. Oil level in the sight glass was low, need to think about an oil change soon.

11 Dec 2020 Friday 09:44 [Greg, Devon] Replaced ion gauge on 65R. Roughed out chamber and beam line.

10 Dec 2020 Thursday 12:10 [Devon Jacobs] Came in this morning to find a phase failure alarm. Found many of the usual suspects, ion gauges, gate valves, in an off state. Of particular note was the alphatross vacuum pumping system. We didn't discover the source vacuum was out until much later. The only indication of such a failure is visual inspection of the vacuum control module in the source area, or if we had thought to look at the software which would have been displaying whatever it would with the ion gauge off. Going forward, should consider tapping into vacuum control circuitry to establish an alarm or other kind of indication for a failure. In other news, facility shutdown is progressing. We've also found the process cooling water emergency dump valve to be inop. Solution in work...

04 Dec 2020 Friday 11:34 [carter, leblanc] Compressor blew fuses again. A new set of contacts has been installed. Also a missing bolt was replaced on the solenoid actuator. Hopefully one or the other was the cause of blowing fuses.

02 Dec 2020 Wednesday 16:08 [carter] Test of addition of Greg LeBlanc to logbook "mail to all operators" function.

02 Dec 2020 Wednesday 16:07 [carter] Test of logbook "mail to Greg LeBlanc" addition.

18 Nov 2020 Wednesday 14:37 [carter] The main air compressor went out again today during the run shutting down the machine. Found one blown fuse and one thermal overload tripped again, same as before. This time I have burnished the contacts as this is a possible cause. I will also check on the availability of a new set of contacts for this contactor. Part number is Square D TA81. This set was replaced for the same symptom April 2006.

18 Nov 2020 Wednesday 11:41 [carter] Currently there are software imposed limits for the alphatross controls of the following, Mass Flow for HE3 and HE4 of 0.18 to 0.50 SCCM and also a software limit on the oven voltage of 90 volts max.

18 Nov 2020 Wednesday 11:25 [Tom Massey] Alphatross start up normal. Starting current 1.1-1.2 microamper. This was slighly unstable. Dr. Ingram saw this behavior also but this eventually stablized at the higher value. His test showed that the source mass flow can be reduced to ~0.18 CCM with improved current output. We currently have 200 nA on target with no current on the collimator at the 65R HeBGB target position.

13 Nov 2020 Friday 11:23 [Devon] The beam right object slit is repaired. Took the micrometer from the Y-up slit. Did not have to break vacuum. The failed beam right slit was not at or beyond the zero point, it seemed to be sitting at around the 0.03" open point. But that is only an estimate as to how far open it was. New Beam right zero set when slits were in contact with beam left at zero. The Y-up slit is trapped in place at > 0.4" setting with a hose clamp to prevent the spring from pushing the slit out further.

11 Nov 2020 Wednesday 15:27 [Tom Massey] The neutron sources have been returned to the source room.

08 Nov 2020 Sunday 10:38 [Tom Massey] The beam right object slit no longer moves. Trying to turn by hand is not possible. The shaft does not move at all. We have been switching the object and image slits between 0.030 and 0.200. This is done to determine the energy by the magnet and calibrate the nmr at the smaller setting of 0.030. The slits are then opened to get enough beam to obtain enough counts above background.

07 Nov 2020 Saturday 07:54 [Tom Massey] The alphatross has 1.1 microamps this morning.

06 Nov 2020 Friday 09:14 [Tom Massey] The swinger turbo pump's roughing pump was making loud noises this morning. The turbo pump has been turned off.

04 Nov 2020 Wednesday 16:16 [carter. shafer] At 11:30 PM last evening the main air compressor stopped. Upon inspection found and repaired the following, blown fuse, thermal protector tripped in motor starter relay box, conduit feed termination had worked loose and needed re-connected at motor and a long bolt (1 of 4) that holds the motor bearing housings together had worked loose and needed replacing. Doug had a new piece (1/4 in. all-thread) in the shop to replace the bolt. After the repairs were made the motor started and the compressor is working again.

03 Nov 2020 Tuesday 23:30 [Doug Soltesz] Tonight at 11:20 pm, the "water hot or air low" alarm sounded and the grid voltage quickly increased to around -15 V. The accelerator was immediately shut off and the magnets turned down. We contacted Devon, who said it was probably the air compressor, and he will be in to fix it on Thursday.

03 Nov 2020 Tuesday 23:30 [Doug Soltesz] Tonight at 11:20 pm, the "water hot or air low" alarm sounded and the grid voltage quickly increased to around -15 V. The accelerator was immediately shut off and the magnets turned down. We contacted Devon, who said it was probably the air compressor, and he will be in to fix it on Thursday.

02 Nov 2020 Monday 08:09 [phase failure detected] The phase failure detected alarm on the E2 panel was on this morning. I checked all vac systems and everything seems to be operating normally. Best guess is there was a very short power outage that set off the alarm but did not hurt anything. I will proceed to turning on the source unless I hear otherwise. -KB

30 Oct 2020 Friday 13:31 [carter] Emergency generator oil changed by facilities.

28 Oct 2020 Wednesday 17:40 [carter] The fictitious chain stretch alarm was caused by sparking at the alphatross source. The cause was a diagnostics induced disconnected plasma box fan cable that was not plugged in and hanging from the source platforming sparking to ground. Initial restart of the source produced 450 nA of beam within 15 minutes.

28 Oct 2020 Wednesday 12:48 [Tom Massey] The chain stretch alarm is going off. This happened when the source was being brought up. This came on when the bias voltage was being raised. The alarm was gone when the bias voltage on the alphatross was reduced. When the extractor voltage reached above ~5 kV the alarm would be set off. Sparking???

23 Oct 2020 Friday 15:44 [Tom Massey] The accelerator was conditioned up to 4.5 MV. We are currently running at 3.371 MV.

15 Oct 2020 Thursday 17:15 [Devon] Alphatross update: All 10 g of Rb loaded without issue. Flowed He thru the source over lunch. Then went for operation. Plasma struck right away but vacuum reading was higher than expected as Probe voltage came up. Let it run a couple hours as suggested in the manual, vacuum came down to expected readings. Our instructions have been changed to say the Probe runs at 6 KV instead of 6.5 KV. Not sure how we got to 6.5, but the manual and other labs use 6. Moved on to oven heating and applied voltages. The focus is a key player in tuning now, which we didnít see when the focus element was shorted by Rb. Initially the beam output was less than 100 nA and the gas flow was higher than expected. As time went on a lower gas flow started giving higher beam output, now closer to the expected .2 sccm. I recall seeing this the last time the bottle was changed. The Rb oven behavior is different now. I had used a bit of thermal contact grease on the TC last time. This time I packed the connection heavier with grease. The oven was set to 87.5 V, what had become our standard and gave us a temperature reading in the upper 180s C. We had seen instability appear if the temp reached 190 C around 90 V. After some time of running today it was noticed that the oven temp had already reached 191 C at 87.5 V. The beam had also come up to 250 nA. Lowered the oven volts to get temp back down and then started nursing it back up. We are currently at 185 C at 77.5 V and getting 200 nA. This makes me think that we were not getting good TC readings on the last build and actually over driving the oven. Something had to make the Rb mess. Iím going to leave the source running overnight as it is right now. Perhaps the beam will improve by morning. If or if not, there should be some exploration of new operating parameters on the oven. No idea what the upper limit is at this point. If somebody wishes to explore this tomorrow then great. If not, please shut down the source. Also, the big einzel lens pretty much does nothing. If you raise it high enough it defocuses the beam a little. Another item for further contemplationÖ

08 Oct 2020 Thursday 11:39 [Devon] Alphatross update: It took many iterations of wiping with Kimwipes and water, acetone, alcohol, and heat gun to get the voltage gap insulators to measure >4000 on the MegOhm meter. The extractor gaps were worse than the acceleration tube. The extractor/lens core components all had to be bead blasted, the Al items were fairly pitted by Rb. The source is back together and under vacuum. New plasma bottle and canal installed. Nothing remarkable about right/left alignment on the probe electrode, but it was observed that pushing on the gas line can move the bottle right/left some. The Ta canal tube had eroded on the exit end, almost thru to the Al base plate. Seems that this is the heaviest wear item along with the Al canal holder button. The new RGA is installed and running. Initial analysis of just vacuum showed no indication of leaks, typical High Vac spectrum according to DCI, although there is a hydrocarbon peak, which is odd since this is a dry pumped system. Now flowing He for clean up and watching the RGA. Masses 28 and 32 come on pretty strong when the gas flowsÖwill continue to monitor this.

29 Sep 2020 Tuesday 11:59 [Devon] Alphatross update: The CEC is clean and re-assembled. Removed the extractor/lens assembly. Puddle of Rb behind (downstream) the extractor electrode opening. Good look down the acceleration tube, there is a big collimator plate in front of the steerers, very large blue stain on the Ta plate which is centered. Very large gold stain on the Ta plate which is off to the right and low. Not sure what to make of the size, color, or location of these stains. Beginning the tear down and cleaning of the extractor/lens...

28 Sep 2020 Monday 15:04 [Tom Massey] All of the neutron souces have been returned to the source room.

25 Sep 2020 Friday 15:38 [Devon] Alphatross notes on disassembly: Iíve posted a drawing of the source on the handbook to help with any references below (why does the Comdoc always make a crooked scan?). Rb neutralized without much fanfare. The greatest reactions were in the vault during disassembly, but not enough to register on the VESDA fire alarm. Care and patience must be taken when removing the oven, the wire wick has lots of Rb that gets re-exposed as the wick is moved. The water tub outside was unremarkable. Not much Rb left in the oven, guessing maybe 20% or less of the initial charge. The plasma bottle has milky white coating between the RF bands which Iím thinking is an expected result. The insulator and exit canal seem to be in quite good shape. Still plan to use a fresh bottle and clean/align the probe assembly (gas end, items 28-36). The Al canal holder, item 20 (exit aperture) was again eroded off axis to the right. This is why I want to take a good look at the probe end of the bottle. The holder was not very caked with Rb. Next we get to the charge exchange chamber (CEC) entry cooling baffle (item 48), its aperture was well occluded, less than a quarter of the original diameter (6.4mm) by eye, and the remaining opening was off axis right. The CEC itself (all of item 17) didnít seem too bad, nothing like the ANSTO pictures on SNEAP. The entry aperture was about half occluded (nominally 4.8mm), and there was some Rb puddling at the base of the flange where it meets the entry cooling baffle. The CEC exit was interesting. The Source Exit Spacer (Stainless steel, item 3) has a much larger hole (12.75mm) and conically expands after the hole until it meets the titanium exit aperture (item 4) which has the same size hole. The idea is that Rb should condense on the exit aperture and flow back to the oven thru the drain hole in the bottom of the conical spacer as part of the Rb cycle. Some Rb had wicked into the contact gap between aperture plate and spacer, but the drain was open, so one would assume the cycle is working. Any Rb that escapes the exit aperture is confronted by the exit cooling baffle (item 52, 19mm diameter) and the large space between the source base plate (item 1) and the baffle plate. This space had a fair amount of Rb pooled in the bottom, the greatest concentration other than the oven. Rb that collects in this space is gone from the cycle. The next components of the source are the extractor/lens (focus) assembly followed by the acceleration tube (bias). Gap resistance was checked on these devices while the source was still under vacuum. The 300 MOhm resistor gaps on the acceleration tube measured 240 MOhm consistently. The 150 MOhm resistor gaps on the extractor measured 100 MOhm consistently. Didnít think I needed to check continuity btween the lens electrode and the adjacent extractor gaps. We re-checked resistivity after the source was vented to air. The acceleration gaps had now dropped to 80 MOhms and decreased continuously towards the CEC. The first extractor gap had dropped to 1 MOhm and the others got less moving towards the CEC. Checked the lens to the adjacent extractor plane closest to the CEC, 300 KOhms; adjacent plane further from the CEC, 600KOhms. Both of those measurements should have been infinite since the lens element is isolated. Just for fun, removed one 150 MOhm resistor, measured 149 MOhm. So, lots more cleaning and disassembly ahead of us.

25 Sep 2020 Friday 15:12 [Tom Massey] The Pu(Be) and 252Cf sources have been placed in the back of the tunnel. Kristyn is doing calibrations and needs a low neutron background.

24 Sep 2020 Thursday 13:59 [Devon] Alphatross notes prior to disassembly: 881.6 hours. Generally suffering from low beam output and instability. After the last power outage and Ar backfill the source sat under vacuum for several weeks of sputter source runs. Upon start up there was immediate plasma ignition with a pinkish hue. Source was left running in hopes of clean up, the pinkish hue faded to aqua quickly. He- was initially on the order of 100-300 nA in the LE cup with instabilities on the order of 5-10 Hz. After several days the output had increased to around 500 nA but the source had to be shutdown for sputter pellet change. On restart the plasma would not ignite, had to raise the gas to .5 sccm and raise the probe to 5 KV to get ignition. After another 4 days the current had increased to 800 nA with instability on the order of 1-2 Hz. During RBS work on day 5 the Probe was increased from 6 to 6.5 KV which gave an increase in beam reaching the RBS chamber. But from that point forward the beam suffered a steady downward slope for a few days until the output had reached 100 nA again, and the magnitude of the instabilities had increased, but not the frequency. Attempts at tuning couldnít produce any more than 150 nA at the peak of the instability cycle. The BPM showed a reasonable looking profile at times, but it was also swamped with baseline noise, the effect seen when ions strike the housing. In some cases the baseline noise would persist for many seconds after a voltage change in the system. Nothing left to do but tear down the sourceÖ

24 Sep 2020 Thursday 12:05 [carter] OIT has returned the portable phone for the vault claiming is is OK. I have returned it to its normal place over the swinger scope inside the vault door. We will see.

21 Sep 2020 Monday 14:53 [Gula Hamad] Note on BCI Calibration: BCI test was done using Both ranges 1E-06 and 1E-07 and we got the same result.

21 Sep 2020 Monday 14:33 [Gula Hamad] BCI Calibration: We did test the BCI by sending a known current and verifying the BCI reading. We calculated the measured charge from BCi = 5.97E-04 [C] and the Applied current = 6.00E-04 [C]. The ratio between charge from BCi to the Applied current = 0.9955

19 Sep 2020 Saturday 13:46 [Ingram] Alphatross note: I saw that the probe was at 6.0 kV. I decided to increase it to 6.5 kV. The beam on target ( I am taking data) increased, it doubled from 10 nA to 20 nA.

18 Sep 2020 Friday 12:59 [carter] With reference to the buncher no output issue, I decided to double check that the tank switch was indeed in position 2, rotated the switch completely around and back to position 2, the position we wanted for alphas. Re checked the operation. Saw a glimmer of 5 MHZ on the output and tuned the output tune control many turns and found the resonance, so it appears that either the tube selector switch was not completely in position or possibly that we don't frequently run in position 2 and it needs to be re-tuned significantly to resonate. Anyway, it appears that the buncher is again OK and tuned for the 16 cm bunch tube. Pulsed beam should now again be possible.

18 Sep 2020 Friday 11:15 [carter] The portable phone that was located in the vault has been picked up by OIT for diagnostics and repair. There is no phone at that location now so use the portable from the control room if phone service in that area is needed until further notice.

18 Sep 2020 Friday 11:10 [carter] The buncher has developed a failure overnight. Restricted access diagnosis is a defective output tube screen power supply. DC operations can continue with the machine as scheduled. Maintenance on the buncher is scheduled for Sept 30 unless an opportunity presents itself and the sources are down before that date.

17 Sep 2020 Thursday 09:12 [Devon] Magnet Primary pump running. Resume schedule.

16 Sep 2020 Wednesday 15:49 [Devon] A viable bearing assembly for the magnet pump was found in the shop. Still has bearings, yay. Hope to have the existing pump up and running tomorrow, for as long as the motor lasts.

16 Sep 2020 Wednesday 15:45 [Devon] Additional videos have been added to the bottom of the handbook page. There is a video of navigating the control room camera station and a startup and shutdown video for each source. Enjoy.

15 Sep 2020 Tuesday 15:55 [Devon] The Magnet Primary cooling pump failed and is unrepairable. Attempted to install a new shaft seal and coupling but wear and pitting were too bad to get a seal. This was only going to be temporary anyway due to severe slop in motor bearings. Checking on price and lead time for a replacement...

15 Sep 2020 Tuesday 11:29 [Devon] Magnet Primary water pump failed. Attempting rebuild...

14 Sep 2020 Monday 16:22 [carter] Capacitor in LE tube vacuum gauge was shorting and was replaced. 0.5 uF 4000 VDC. Last spare on the shelf but have a complete spare gauge power supply unit that was removed from old duoplasmatron source as a supply of spares.

11 Sep 2020 Friday 11:51 [carter] Analyzing magnet power supply would not run up. Problem traced to the collector mount screw on Q-7 that had loosened, probably because of thermal cycling. Will add star locking washers to these screws to help prevent this.

05 Sep 2020 Saturday 14:39 [Ingram] pellet changed to 12C. Left source gate valve closed.

04 Sep 2020 Friday 16:43 [Devon] The zero gauge reading was on the suction side of the comfort chilled water pump. The make up regulator was stuck and there was a leak in the Trane chiller. The regulator set point was exorcised and is working again. The leak was a small valve with a cracked body feeding an air bleeder inside the chiller. Closing the valve stopped the leak for now. There is a work order to replace the cracked valve.

04 Sep 2020 Friday 11:23 [Ingram] Beam instability back so that I have switched to GVM control. This coincided with instabilty in the extractor current in the alphtross.

04 Sep 2020 Friday 10:40 [Ingram] Water in the mechcanical room above the control room is coming from three sources. Water from the HVAC air handler, as usual. Water from a roof leak probably from the rain yeasterday giving a pool of water around the water heater. This is dripping very slowly through to the control room and onto the table close to my laptop, that is why I noticed. The third source is in the inner room around the water pump. The blue cylinder connected to the water pump is sweating alot. Also one of the water gauges connected to the HVAC pump is reading zero.

04 Sep 2020 Friday 10:36 [Ingram] Alphatross producing about 0.7 uA LE cup. Stable. Running 10 nA on target and taking RBS and PIXE data. Things of note about the alphatross, to me, Einzel lens, Focus Voltage and both steerers are at zero volts. They either have no effect on the beam or in the case of the steerers, reduce the current. No obvious reason why the beam current has grown over time or why the beam is now more stable.

03 Sep 2020 Thursday 12:04 [carter] Replaced V6 6AH6 120 Hz drive tube in the NMR to restore the full horizontal sweep in the 120 Hz position. Adjusted the sweep gain to match the 60 Hz.

02 Sep 2020 Wednesday 15:47 [carter] Created a wtchdog timer for the accelerator logging system. If the logging fails to update the tandem.time file over a 5 minute period, the system will automatically restart the acelerator logging program.

02 Sep 2020 Wednesday 08:20 [carter] Looks like accelerator logging system stopped at 2:23:54 AM this morning, possibly power glitch that shut off power to sputter source HV platform? Anyway, restarted the logging system at 8:12 this morning. Looks like it is working OK now.

31 Aug 2020 Monday 08:39 [Devon] The front door is being replaced this morning.

28 Aug 2020 Friday 12:18 [carter] The alphatross computer has completed the new disc image. The new image was installed as the current system disk, the old one retired as a backup disc. The new system has been updated to the current centos7 version 7.8.2003. Tandemwatch and inflection controls restored to normal.

27 Aug 2020 Thursday 12:41 [carter] I have begun an image backup of the alphatross computer prior to applying system updates for Centos7. The inflection magnet control is available on the tandem computer and the tandemwatch (downcharge monitor program) is displaying on the congrol room ddaq computer. I will restore the alphatross computer to normal operation Monday AM.

24 Aug 2020 Monday 15:20 [Zach Meisel] Operators needed (week of September 7): Greetings operators, Alexander will be running 12C on Mg targets in the charged-particle time-of-flight chamber on R65 to study the statistical properties of nuclides in this region. He would like operator help to do this. If possible, we will try for overnight runs. These runs will mostly involve sitting at a single energy and letting the data come in. Please fill-out this Doodle poll to indicate your availability: https://doodle.com/poll/uwrhze3nnkada3sb. Thanks for your help! -Zach

24 Aug 2020 Monday 12:52 [Cody Parker] Came in to stop the overnight TexAT calibration and the SF6 alarm was going off. I lowered the set point slightly and the alarm stopped. I checked the tank pressure and temperature, and both have dropped some. I talked to Don on his way to the LTR and he will monitor the plots.

23 Aug 2020 Sunday 20:25 [Cody Parker/Jack Bishop] We have completed the 12C(n,n2) and 16O(n,a) runs using TexAT! Thank you to everyone for all of your help! We are working on a calibration source run with TexAT overnight, and plan to start the ~week-long process of disassembly and packing tomorrow. The D2 gas cell is still pressurized and the water cooling is on. All other accelerator-related equipment and the sputter source have been shut down.

23 Aug 2020 Sunday 00:20 [Devon] VESDA minor smoke event in the small target room. Definite smell of something burned, but no obvious source, detection returned to baseline. Will have to be on the lookout for something not working in there. Came in to reset the fire alarm, experiment resuming.

12 Aug 2020 Wednesday 14:07 [carter] Trane replaced other temperature sensor today and both circuits are said to be OK and running now.

11 Aug 2020 Tuesday 16:14 [Devon] The gas is in and drying. The accelerator is ready to run tomorrow as long as the AC stays on.

11 Aug 2020 Tuesday 11:49 [Devon] Chiller update: A factory tech was here this morning. He happened to have a water temperature sensor. It has been installed and we should be back to half a chiller. They identified another bad sensor that they did not have on hand. Will be back tomorrow late morning or early afternoon to install the other sensor. People will have to be on the look out for the factory tech, and pause the run or survey the roof, and let him in the roof gate to make the repairs. I have started to evacuate the accelerator tank. Gas transfer will be done by 5 pm if the door seal doesn't leak.

11 Aug 2020 Tuesday 07:17 [Devon] The Facilities work center page is still not functioning on VPN. The issue was reported to OIT yesterday morning. There was an urgent work order for the chiller in early July that was never addressed. I can't obtain that work order number until I get to campus.

11 Aug 2020 Tuesday 07:13 [Devon] So the chiller was completely dead by around 7 pm last night. First person in the lab today needs to call Facilities and report a chiller emergency, and stand by to let in the technician. I won't be there until 9:45ish. From what I discussed with the technician yesterday, a reliable cure for this chiller is not certain and certainly not timely. The thing is a lemon. Management from both organizations need to discuss a timeline, consider installation of a portable, temporary chiller, and discuss the disposition of our out-of-state visiting researchers.

10 Aug 2020 Monday 14:48 [Devon] Facilities Management was here most of the day. Both halves of the chiller are still running at this moment. There is still a problem such that the measured water temperature in the unit and the actual water temperature are different. They will order new sensors. But its running. Plan is to transfer gas tomorrow after operator class.

10 Aug 2020 Monday 11:53 [Devon] Tank Opening: 7173 hours. We observed a sudden onset of noise in the LE Tube downcharge while running at 2.7 MV. We entered the tank expecting to find an open resistor or broken connecting spring. We found neither. We started the process of looking back at logged data. The onset of the noise was related to the tube entrance lens being driven to a voltage above the terminal induced voltage of the downcharge on plane 4 where the lens is. We investigated the performance of just the lens with vacuum in the tube but 1 atm air in the tank. The lens voltage and current were stable and giving the expected values. There is a spark gap on the lens feed path. The points of the gap were dirty and sparked at around 8 KV. We insulated the gap with Teflon sheet and ran at around 15 KV without instability, confirming that the supply and wiring seemed to be ok. A review of the prior month data showed that the noise had been intermittently present when the lens was driven above the induced voltage, or when the lens had been left on while the accelerator is off (donít leave electrostatic devices on!). Given the compulsion to do SOMETHING while the tank was open, decided to clean the spark gap with sandpaper and kimwipes. Nothing else to do with respect to this problem. We will have to monitor the noise behavior going forward. It had also been a year since the bearing disaster. A year is the target timeframe for greasing bearings, so that was done. No other maintenance due to schedule pressure.

10 Aug 2020 Monday 06:34 [carter] Same thing for me, says not on VPN but I am also.

10 Aug 2020 Monday 06:17 [Devon] VPN test. Facilities web page says I'm not on VPN, page won't work. VPN says connected, this page works...

07 Aug 2020 Friday 09:48 [Devon] During the monthly safety inspection, found the vault phone to be severely locked up. Had to remove the battery. Seems to be OK now. Radiation?

05 Aug 2020 Wednesday 16:26 [Tom Massey] We were running at 2.5 MV on the terminal today. Devon noticed that the LE current has been jumpy. This did not change when the LE cup was put in. These excursions seemed to reduce in size with time. The changes were most visible on the Tandem watch screen.